The manufacturing overhead budget at Cutchin Corporation is based on budgeted direct labor-hours. The direct labor budget indicates that 3,500 direct labor-hours will be required in September. The variable overhead rate is $7 per direct labor-hour. The company's budgeted fixed manufacturing overhead is $43,190 per month, which includes depreciation of $3,570. All other fixed manufacturing overhead costs represent current cash flows. The September cash disbursements for manufacturing overhead on the manufacturing overhead budget should be:
Answer #1
Concepts and reason

Manufacturing Overheads: Manufacturing Overheads are the aggregate of Indirect Material, Indirect Labor and Indirect Expenses which are incurred during Production of Finished Goods in the Factory. These are the cost which cannot directly belong to a particular unit of a product.  

Fundamentals

Manufacturing overhead rate: It is a predetermined rate for overheads incurred during the manufacturing process. It is computed by dividing the total estimated amount of manufacturing overheads with an estimated value of the allocation base.

The basis of allocation can cost of direct material, direct labor, the output produced, direct labor hours or machine hours. The basis of allocation is decided on the basis of cost behavior.

Variable overheads: It is the expense directly related to the product. The expenses which can be directly traceable to the product are included in the variable expenses. The units sold during the period are multiplied by the variable cost per unit in order to determine the amount of variable expenses.

Fixed overheads: The fixed expenses are those which do not change with the level of activity. If the company does not produce a single unit, then also the occurrence of these expenses are fixed. When fixed expenses can be classified between different lines of product, different departments on the specified basis, then those fixed expenses are called traceable fixed expenses.

Compute variable manufacturing overheads of Company C as given below:

Variablemanufacturingoverheads=Variablemanufacturingoverheadrateperdirectlaborhour×Directlaborhours=$7×3,500hours=$24,500\begin{array}{c}\\{\rm{Variable manufacturing overheads}} = {\rm{Variable manufacturing overhead rate}}\\\\{\rm{ per direct labor hour}} \times {\rm{Direct labor hours}}\\\\{\rm{ = \$ 7}} \times {\rm{3,500hours}}\\\\{\rm{ = \$ 24,500}}\\\end{array}

Compute the cash payments during the month of September as given below:

Cashpayments=BudgetedfixedmanufacturingoverheadsDepreciation+Variablemanufacturingoverheads=$43,190$3,570+$24,500=$39,620+$24,500=$64,120\begin{array}{c}\\{\rm{Cash payments}} = {\rm{Budgeted fixed manufacturing overheads}} - {\rm{Depreciation}}\\\\ + {\rm{Variable manufacturing overheads}}\\\\{\rm{ = \$ 43,190}} - \$ 3,570 + \$ 24,500\\\\ = \$ 39,620 + \$ 24,500\\\\ = \$ 64,120\\\end{array}

Ans:

Cash payments during the month of September are $64,120.